hi i have been trying to do these equations for such a long time and i just cant do it please help me
Write an equation of the line that is parallel to the given line and passes though the given point.

10. y= -3x+2 (2,3) 11. y=1/2x - 5 (-3,-1)

Answer by scott8148(6628)

lines that are parallel have the same slope (the m in y=mx+b), but have different y-intercepts (the b term)

when a line goes through a point, it means that the x and y values of the point satisfy (fit) the equation of the line

y=-3x+b for x=2, y=3 ... 3=-3(2)+b ... 3=-6+b ... b=9 ... y=-3x+9

y=(1/2)x+b for x=-3, y=-1 ... -1=(1/2)(-3)+b ... -1=-(3/2)+b ... b=1/2 ... y=(1/2)x+(1/2)